OpenStudy (anonymous):

I NEED HELP ASAP!!! PLEASE?!?!?!? Jose received $30 for his birthday that he put in his piggy bank. Each week, he puts 4 more dollars in his bank so that the amount of money in dollars D in his bank is given by D = 4w + 30 where w is the number of weeks. His sister Marisol makes a table of reasonable values for w to determine how much money Andre will eventually have. If one of the values for w Marisol selects is 16, what would be her corresponding entry for D? (Points : 1)

OpenStudy (anonymous):

@BunnyBree Can you help me?

OpenStudy (anonymous):

@mclaughlin23

OpenStudy (anonymous):

@Lyrae

OpenStudy (anonymous):

@Sparklestaraa

OpenStudy (sparklestaraa):

OK let me attempt this :D

OpenStudy (sparklestaraa):

D=4w+30 One entry= 16 4*16= 64 D=64+30 D=94 @iLorita

OpenStudy (anonymous):

Thank so much!! @Sparklestaraa

OpenStudy (sparklestaraa):

Your welcome :D Wish you the best if u need anymore help just tag me in :D

OpenStudy (anonymous):

Melania has two jobs. During the day, she works as an office clerk, and in the evening she works as a cashier. Her office job pays her $13 per hour, and her cashier job pays her $9.25 per hour. In one week, Melania worked 38 hours, and she earned a total of $434. If c represents the number of hours that Melania works as an office clerk, write an equation that can be used to find how many hours Melania worked in both jobs. @Sparklestaraa

OpenStudy (sparklestaraa):

$434= $13c+$9.25h h=hours worked as a cashier c- hours melania worked in both jobs $434= total price $13= Office jobs payment per hour $9.25=Cashier Jobs Payment per hour.

OpenStudy (anonymous):

It doesn't have that in the options :/

OpenStudy (anonymous):

13c + 9.25(c – 38) = 434 13c + 9.25(38 – c) = 434 13c + 9.25(434 – c) = 38 9.25c + 13(38 – c) = 434

OpenStudy (sparklestaraa):

write that again but this time don't copy and paste :P so i can read it or change the little signs in it or screenshot and upload :D

OpenStudy (anonymous):

A: 13c+9.25(c-38)= 434 B: 13c+9.25(38-c)= 434 C: 13c+9.25( 434-c)= 38 D: 9.25c+13(38-c)= 434

OpenStudy (sparklestaraa):

well first if u read it says How many hours in one week all together which is 38 and we r not trying to use price at all so U can single it out to just C :D

OpenStudy (anonymous):

Oh thanks!! lol I feel so stupi xD

OpenStudy (anonymous):

Can you check if I did this correctly? I promise this is the last time I ask you :)

OpenStudy (sparklestaraa):

don't feel stupid xD be happy I'm here :P Its fine i will check I'm in 9th grade i do the same so i can totally understand :D

OpenStudy (anonymous):

1. What algebraic property does this statement show? 3 + (7 + 2) = (3 + 7) + 2 associative property<< closure property commutative property symmetric property Question 2.2. Solve. 2x – 7 = 6x – 15 2<< 3 5 6 Question 3.3. What is the slope of the line that passes through the points (1, 3) and (5, –2)? A: 5/4 B: -5/4<< C: 4/5 D: -4/5 Question 4.4. The sum of three consecutive integers is –27. What are the numbers? {–5, –9, –13} {–10, –9, –8} {–7, –9, –11} {–11, –13, –15} Question 5.5. Which equation can be used to solve the following word problem? Jessica has 9 more nickels than dimes, and the total value of her coins is $1.50. How many nickels does she have? Let n represent the number of nickels. 0.10n + 0.05n + 9 = 1.50 0.10(n – 9) + 0.05n = 1.50<< 0.10n + 0.05(n + 9) = 1.50 0.10(n + 9) + 0.05n = 1.50 Question 6.6. If A = {2, 4, 6, 8, 10} and B = {1, 2, 3, 4, 5}, find A 09_04_ST_15.gif B. {1, 2, 3, 4, 5, 6, 8, 10} {1, 3, 5} {2, 4}<< {7, 9} Question 7.7. Select the graph that represents the inequality. 2x + 3y ≤ 6 A: http://static.k12.com/eli/ecollege/543/9_29484/2_31257_4_29487/3069121b90f3da6bf2044f3bf82e2f3f2df8ce93/media/622039939f96f1a9a39f2845ecd38f4791d3b476/mediaasset_212708_1.gif B: http://static.k12.com/eli/ecollege/543/9_29484/2_31257_4_29487/3069121b90f3da6bf2044f3bf82e2f3f2df8ce93/media/b062168b35202e07967747a76c0d83d458a26234/mediaasset_212709_1.gif C: http://static.k12.com/eli/ecollege/543/9_29484/2_31257_4_29487/3069121b90f3da6bf2044f3bf82e2f3f2df8ce93/media/0defeb5d31667f6c849b6c88eb380e700a04b51d/mediaasset_212710_1.gif D: http://static.k12.com/eli/ecollege/543/9_29484/2_31257_4_29487/3069121b90f3da6bf2044f3bf82e2f3f2df8ce93/media/95bb7389e48763a4b4e3b3d2e27e5c5c4609fb0a/mediaasset_212711_1.gif Question 8.8. Select the pair of equations whose graphs are perpendicular. A: 5x – 8y = 9 12x – 5y = 7 B: 3x + 6y = 8 y = 3x – 8 C: y = 2x – 7 x + 3y = 3 D: 2y = –3x + 5<<< 2x – 3y = 4 Question 9.9. Combine like terms to simplify the expression: 7x + 3y – 2 + 6x – 1 + y2 2y3 + 13x – 3 3y3 + 13x2 – 3 y2 + 3y + 13x – 3<< y2 + 3y + 13x2 – 3 Question 10.10. Which is an equation in slope-intercept form for the line passing through (–1, 5) and (5, 8). A: -1/2X+ 11/2 B: -1/2X- 11/2 C: I/2X- 11/2 D: 1/2X + 11/2 Question 11.11. Given the linear equation x – y = 5, find the slope of its graph. –5 –1 1<< 5 Question 12.12. Given the linear equation 3y + 4x =12, find the y-intercept. A: -4/3 B: -4 C: 4/3 D: 4 Question 13.13. Evaluate the expression when a = –1.4 and b = –2.7. |a| – b 4.1 –1.3<< –4.1 1.3 Question 14.14. Solve for x. X-1/12= 7/2 x = 2 x = 6 x = 41 x = 43 Question 15.15. Graph the solution of the inequality. 14 ≤ 2t + 18 < 20 A: http://static.k12.com/eli/ecollege/543/9_29484/2_31257_4_29487/3069121b90f3da6bf2044f3bf82e2f3f2df8ce93/media/792123b1c375a1030f5bf800d19afbd942c2d406/mediaasset_212712_1.gif B: http://static.k12.com/eli/ecollege/543/9_29484/2_31257_4_29487/3069121b90f3da6bf2044f3bf82e2f3f2df8ce93/media/4d19f86e581d86c301d85c72301bc7fd035007b8/mediaasset_212713_1.gif C: http://static.k12.com/eli/ecollege/543/9_29484/2_31257_4_29487/3069121b90f3da6bf2044f3bf82e2f3f2df8ce93/media/29de64657257e1042cdfd7dd082c2db69364c8df/mediaasset_212715_1.gif D: http://static.k12.com/eli/ecollege/543/9_29484/2_31257_4_29487/3069121b90f3da6bf2044f3bf82e2f3f2df8ce93/media/e6f7674f86d21dca774a96d1ef7bf7d52f1a712e/mediaasset_212716_1.gif Question 16.16. Solve the system of equations. 3x + 3y = –6 –x + y = 0 x = –1; y = –1<< x = 10; y = –3 x = 22; y = 7 x = 14; y = 5

OpenStudy (anonymous):

Sorry I toke soo long it was a lot of typing lol

OpenStudy (anonymous):

the ones that have << are the ones that I picked as an answer

OpenStudy (anonymous):

@Sparklestaraa

OpenStudy (sparklestaraa):

lol u still can't see a lot try to screenshot it

OpenStudy (sparklestaraa):

I'm reading it sorry i went to eat @ilorita

OpenStudy (sparklestaraa):

Module 1- Lesson what xD

OpenStudy (anonymous):

I can't I'm in my computer lol

OpenStudy (anonymous):

@Sparklestaraa Sorry I didn't reply I went out to eat as well

OpenStudy (anonymous):

Which one are you talking about? @Sparklestaraa

OpenStudy (anonymous):

That's for number 7

OpenStudy (anonymous):

tHAT'S # 15 @Sparklestaraa

OpenStudy (anonymous):

Can you please check them @Nnesha

OpenStudy (anonymous):

can you help me?tag @DanJS

OpenStudy (sparklestaraa):

The tag is from me too :D @DanJS and hey @dtan5457 can u spare some help :D

OpenStudy (sparklestaraa):

@pooja195

OpenStudy (anonymous):

@hockeychick23

OpenStudy (arabpride):

Please ask one question per post. and Please close this post if you're done. Thanks! :D